OpenStudy (anonymous):

What is the focus of the parabola given by the equation x = 3(y – 1)^2 + 2?

OpenStudy (anonymous):

I keep getting (1, 2&1/12) but its not in my answer choices?

OpenStudy (anonymous):

get that equation into this form: \(\large 4p(x-h)=(y-k)^2 \) since this is a parabola that opens to the right, you must add "p" to the x-coordinate of the vertex to the parabola to get the focus.

OpenStudy (anonymous):

\(\large x=3(y-1)^2+2 \) \(\large x-2=3(y-1)^2 \) \(\large \frac{1}{3}(x-2)=(y-1)^2 \) so \(\large 4a=\frac{1}{3}\rightarrow a=\frac{1}{12} \) so all you do is add \(\frac{1}{12} \) to the x-coordinate of the vertex